# RegMean++: Enhancing Effectiveness and Generalization of Regression Mean for Model Merging

### v1

@@ 1 Introduction
As pretrain-finetune paradigm becomes the foundation of modern machine learning, the number of pre-trained and fine-tuned task-specific models (candidate models) is growing at an unprecedented pace. Model merging (Matena and Raffel 2022; Wortsman et al. 2022; Ilharco et al. 2022; Jin et al. 2022; Yadav et al. 2023; Yang et al. 2024b, a; Yadav et al. 2024), an emerged approach that aims to combine multiple candidate models into a single unified model (merge model) with multi-tasking capabilities, without incurring the computational overhead of traditional multi-task learning (MTL) or full access to original training data.
Regression Mean (RegMean; Jin et al. (2022)), an explainable and computationally efficient model merging method, formulates weight fusion as a closed-form regression problem that minimizes the differences between the outputs of merge model and those of each candidate model. RegMean leverages the inner-product matrices of the input features at each linear layer, including those within the MLP components and attention heads of transformer layers in the candidates. Additionally, RegMean decreases non-diagonal entries in these matrices to enhance stability during merging. The other types of transformer layers’ weights are merged by simply averaging across candidate models. RegMean offers several benefits, including privacy-preserving, computational efficiency, and model-agnostic. These advantages make RegMean one of the most practical merging methods.
However, we discuss an important caveat of RegMean is that it operates by independently applying the closed-form solution to linear layers of the transformer layer, fundamentally ignoring how features and information are processed and propagated through layers in the merge model. These intra- and cross-layer dependencies are crucial for maintaining good representations. The second drawback of RegMean is its memory and storage overhead. RegMean requires memory to store the inner-product matrices, which—for the large-scale models—can rival the number of parameters. In this paper, we aim to address these limitations:
(1) We introduce RegMean++, an alternative to RegMean, which incorporates both intra- and cross-layer dependencies of the merge model’s layers into the RegMean merging objective. RegMean++ consistently improves in ID performance, OOD generalization, and demonstrates sustainability to sequential merging or large-scale tasks. Moreover, RegMean++ shows stronger robustness under various types of distribution shifts with reduced representation bias.
@@ 2 Background and Related Work
Model Soups (or parameter averaging) (Wortsman et al. 2022; Choshen et al. 2022), a well-known approach that performs merging by simply taking the average of all candidate models’ parameters, for enhancing distribution robustness (Wortsman et al. 2022), creating merge models with multi-task or multi-modality capabilities (Sung et al. 2023; Ilharco et al. 2022; Yadav et al. 2023).
Task Arithmetic (Ilharco et al. 2022) introduces “task vector”, a new concept that quantifies the task-specific knowledge of a candidate model, by measuring the difference between the candidate model parameters and the base model parameters. However, Yadav et al. (2023) point out that Task Arithmetic suffers from conflicts where different task vectors update the same parameters in opposite directions. They proposed TIES-Mering, a three-step method that first removes low-magnitude (noisy) parameters, then resolves sign conflicts, and finally aggregates only the non-conflicting parameters. DARE (Yu et al. 2023) proposes to pre-process each task vector independently by applying a Bernoulli mask to randomly zero out its parameters, effectively performing a dropout-like pruning before merging. TSV-M (Gargiulo et al. 2025) extracts task singular vectors via singular value decomposition (SVD) on layer-wise weight updates, selects top-$k$ components for a low-rank representation, and applies whitening to decorrelate subspaces before merging. Marczak et al. (2025) argue that effective merging depends on how well merge updates span each task’s principal subspaces. They propose Iso-C, which sums task-specific updates, performs SVD, and replaces singular values with their average to create an isotropic spectrum. Iso-CTS, an Iso-C’s variant, adds top singular directions from each task’s residual, then orthogonalizes and isotropically scales the result.
Inspired by test-time adaptation schemes (Wang et al. 2020; Liang, He, and Tan 2025), AdaMerging (Yang et al. 2024b) adaptively learns the merging coefficient for each layer of each task vector by minimizing the entropy on unlabeled test data, using it as a surrogate objective to refine the merge model’s performance across multiple tasks. DOGE AM (Wei et al. 2025) applies adaptive projective gradient descent at test time by jointly tuning a small modification vector and layer-wise merging coefficients to minimize prediction entropy on unlabeled inputs. DOGE AM projects updates orthogonally to a shared subspace to resolve task conflicts without hurting the shared knowledge.
Other approaches rely on data statistics, such as Fisher Merging (Matena and Raffel 2022) that requires computing Fisher information matrices. RegMean (Jin et al. 2022) gets rid of expensive gradient computation, formulates merging as a regression problem, then comes up with a computationally efficient and explainable closed-form solution.
@@
We denote matrices by boldface uppercase letters (e.g., $\bm{X}$, $\bm{W}$, $\bm{\Lambda}$), scalars by lowercase letters (e.g., $\alpha$, $l$). For operators, we denote $||\cdot||$ the Euclidean norm and $\text{tr}(\cdot)$ the trace of a matrix. Following  Jin et al. (2022), we consider the training-free merging framework. In such scenarios, we have access to a pool of multiple candidate models. Each candidate model is denoted as $f_{i}:\mathbb{R}^{N_{i}\times d}\to\mathbb{R}^{N_{i}\times|C|}$, and is fine-tuned on a task-specific dataset $\mathcal{D}_{i}=\{(\bm{X}_{i},\bm{Y}_{i})\}$. Here, $\bm{X}_{i}\in\mathbb{R}^{N_{i}\times d}$ is the batch-input of $N_{i}$ samples each of dimensionality $d$; and $\bm{Y}_{i}\in\mathbb{R}^{N_{i}\times|C|}$ is the corresponding target outputs, where $C$ is the set of classes. Our target is to find a merging function that takes a set of $K$ candidate models $f_{i}$, $i\in[1..K]$ and some data points, e.g., the task-specific training samples or held-out out-of-domain samples, as the inputs and returns a merge model $f_{{M}}$. We assume that the model architecture for all models is the same.
Regression Mean (RegMean;  (Jin et al. 2022))
Figure 1: Comparison between RegMean and RegMean++ for model merging. RegMean++ leverages representations from the merge model for merging, enabling accurate alignment with its behavior.
@@
[…]
where $\bm{W}^{(l)}_{M}$ is the merge linear layer’s weights at the same position as $\bm{W}^{(l)}_{i}$ in the model $f_{i}$, $\bm{\Lambda}^{(l)}_{i}=\frac{1-\alpha}{\alpha}\text{diag}\left((\bm{X}^{(l)}_{i})^{\top}\bm{X}^{(l)}_{i}\right)$ is the diagonal matrix, where $0\leq\alpha\leq 1$ is a predetermined scaling factor. Minimizing the loss $\mathcal{L}^{\text{RegMean}}$ in Eqn. 1 means finding $\bm{W}^{(l)}_{M}$ that ➀ approximates the behaviors of all candidate models while ➁ enforcing a regularization that keeps linear layer’s weights of the merge model $\bm{W}^{(l)}_{M}$ close to those of the candidate model $\bm{W}^{(l)}_{i}$. Minimizing $\mathcal{L}^{\text{RegMean}}$ describes a linear regression problem, where the inputs are $\bm{X}^{(l)}_{1},...,\bm{X}^{(l)}_{K}$] and the target outputs are $(\bm{W}^{(l)}_{1})^{\top}\bm{X}^{(l)}_{1},...,(\bm{W}^{(l)}_{K})^{\top}\bm{X}^{(l)}_{K}$]. This objective has a closed-form solution given by:
@@
where $\widehat{\bm{G}}^{(l)}_{i}=\alpha\bm{G}^{(l)}_{i}+(1-\alpha)\bm{\Lambda}^{(l)}_{i}=\alpha(\bm{X}^{(l)}_{i})^{\top}\bm{X}^{(l)}_{i}+(1-\alpha)\bm{\Lambda}^{(l)}_{i}$. Proof can be found in Jin et al. (2022) Appendix A. Other types of parameters in the transformer layer are merged using simple averaging.
@@ 3.1 Motivation
Our first start is to revisit the underlying RegMean’s merging mechanism. RegMean operates by independently applying its closed-form solution to linear layers, including those within MLPs and attention heads, across all candidate models. More concretely, let us fine-grain denote $\bm{X}^{(l,j)}_{i}$ be the input features of the $j$-th linear layer $\bm{W}^{(l,j)}_{i}$ in model $f_{i}$ at transformer layer $l$. RegMean computes the $j$-th merge linear layer weights $\bm{W}^{(l,j)}_{M}$ by Eqn. 2. In this closed-form solution, the merge weight $\bm{W}^{(l,j)}_{M}$ is determined by the individual weights and data statistics $\bm{G}^{(l,j)}_{i}=(\bm{X}_{i}^{(l,j)})^{\top}\bm{X}_{i}^{(l,j)}$, which captures the dependencies among input features across all candidate models. However, this procedure overlooks the information flow and feature transformations that occur both within each layer and across layers of the merge model—dynamics that are crucial for maintaining the merge model’s utilities and generalization.
@@
3.2 Algorithm
We introduce RegMean++, a simple yet effective alternative to RegMean, that incorporates the notion of intra- and cross-layer dependencies between layers of the merge model into RegMean’s objective. Comparison between RegMean and RegMean++ is described in Figure 1. Similar to RegMean, given a $j$-th linear layer at the transformer layer $l$, RegMean++ computes the inner product matrix as $\bm{G}_{i}^{(l,j)}=(\bm{X}^{(l,j)}_{i})^{\top}\bm{X}^{(l,j)}_{i}$. The key difference between RegMean++ and RegMean lies in how input feature $\bm{X}^{(l,j)}_{i}$ is obtained: For input features that are activations (cushion representations between transformer layers), RegMean++ computes $\bm{X}^{(l,j)}_{i}$ based on the activations produced by the previous merge layer $f_{M}^{(l-1)}$ in the merge model, that is, $\bm{X}^{(l)}_{i}=f_{M}^{(l-1)}(\bm{X}^{(l-1)}_{i})$ while RegMean relies on the activations produced by the previous candidate layer $f_{i}^{(l-1)}$ in the candidate model, that is, $\bm{X}^{(l)}_{i}=f_{i}^{(l-1)}(\bm{X}^{(l-1)}_{i})$. All other parameters in the transformer layer, such as embeddings and biases, are merged via averaging. RegMean++ inherits RegMean’s advantages, but introduces a trade-off between model performance and computational cost. During the statistic‐collection phase, RegMean++ incurs additional forward passes, which are computed in the merge model to collect the inner-product matrices, yet the merging time equals that of RegMean. Our RegMean++ pseudocode is described in Algorithm 1.
@@ 4 Experiments
Models and datasets.
Following prior works (Ilharco et al. 2022; Yang et al. 2024b; Wei et al. 2025), we evaluate the effectiveness of multi-task model merging methods on eight standard image classification datasets including: SUN397 (Xiao et al. 2016), Stanford Cars (Cars) (Krause et al. 2013), RESISC45 (Cheng, Han, and Lu 2017), EuroSAT (Helber et al. 2019), SVHN (Netzer et al. 2011), GTSRB (Stallkamp et al. 2011), MNIST (LeCun et al. 1998), and DTD (Cimpoi et al. 2014).
Following Wang et al. (2024), we employ $12$ additional tasks for measuring the sustainability of merging methods. Those include: Flowers102 (Nilsback and Zisserman 2008), PCAM (Veeling et al. 2018), FER2013 (Goodfellow et al. 2013), OxfordIIITPet (Parkhi et al. 2012), STL10 (Coates, Ng, and Lee 2011), CIFAR100 (Krizhevsky et al. 2009), CIFAR10 (Krizhevsky et al. 2009), Food101 (Bossard, Guillaumin, and Van Gool 2014), FashionMNIST (Xiao, Rasul, and Vollgraf 2017), EMNIST (Cohen et al. 2017), KMNIST (Clanuwat et al. 2018), and RenderedSST2 (Socher et al. 2013b).
We assess the performance of merging methods on three CLIP model variants (Radford et al. 2021) with ViT-B/32, ViT-B/16, and ViT-L/14. For the candidate models, we employ off-the-shelf checkpoints from Tang et al. (2024a). Details of these datasets’ description and models can be found in Appendix A.1 and Appendix A.2, respectively.
Comparison methods.
We compare RegMean++ against $11$ recent advanced model merging methods spanning three categories: (1) Data-Free methods: Model Soups (Wortsman et al. 2022), Task Arithmetic (Ilharco et al. 2022), TIES-Merging (Yadav et al. 2023), TSV-M (Gargiulo et al. 2025), DOGE TA (Wei et al. 2025), Iso-C and Iso-CTS (Marczak et al. 2025). (2) Training-Free methods: Fisher Merging (Matena and Raffel 2022) and RegMean (Jin et al. 2022). (3) Test-Time Adaptation: AdaMerging (Yang et al. 2024b) and DOGE AM (Wei et al. 2025). In addition, we also consider the MTL as an upper-bound performance. Details about these methods can be found in Appendix A.3.
Experimental setup.
We employ FusionBench (Tang et al. 2024a) for merging evaluation. We report the accuracy and normalized accuracy (See Appendix B.1 for details) over the test split of all tasks. Hyperparameters for settings are specified in their respective subsections, and further detailed in Appendix B.2. Due to space constraints, we present key results in the main text and defer the full experimental setups and additional results to Appendix B.3 and Appendix C, respectively. Our implementation and guidelines for reproducing those results are attached to the supplemental materials.

### v2

@@ 1 Introduction
As pretrain-finetune paradigm becomes the foundation of modern machine learning, the number of pre-trained and fine-tuned task-specific models (candidate models) is growing at an unprecedented pace. Model merging (matena2022mergingmodelsfisherweightedaveraging; wortsman2022modelsoupsaveragingweights; ilharco2023editingmodelstaskarithmetic; jin2025datalessknowledgefusionmerging; yadav2023tiesmergingresolvinginterferencemerging; yang2024adamergingadaptivemodelmerging; yang2024representationsurgerymultitaskmodel; yadav2024matters), an emerged approach that aims to combine multiple candidate models into a single unified model (merge model) with multi-tasking capabilities, without incurring the computational overhead of traditional multi-task learning (MTL) or full access to original training data.
Regression Mean (RegMean; jin2025datalessknowledgefusionmerging), an explainable and computationally efficient model merging method, formulates weight fusion as a closed-form regression problem that minimizes the differences between the outputs of merge model and those of each candidate model. RegMean leverages the inner-product matrices of the input features at each linear layer, including those within the MLP components and attention heads of transformer layers in the candidates. Additionally, RegMean decreases non-diagonal entries in these matrices to enhance stability during merging. The other types of transformer layers’ weights are merged by simply averaging across candidate models. RegMean offers several benefits, including privacy-preserving, computational efficiency, and model-agnostic. These advantages make RegMean one of the most practical merging methods.
However, we discuss an important caveat of RegMean is that it operates by independently applying the closed-form solution to linear layers of the transformer layer, fundamentally ignoring how features and information are processed and propagated through layers in the merge model, preventing it from generalizing well. These intra- and cross-layer dependencies are crucial for maintaining good representations that influence the final predictions of the merge model.
In this paper, we make the following contributions:
(1) We introduce RegMean++, a generalized extension of RegMean applicable to both vision and language tasks. RegMean++ incorporates both intra- and cross-layer dependencies of the merge model’s layers into the RegMean merging objective. RegMean++ consistently improves in ID performance, OOD generalization, and demonstrates sustainability to sequential merging or large-scale tasks. Moreover, RegMean++ shows stronger robustness under various types of distribution shifts with reduced representation bias.
@@
Figure 1: Comparison between RegMean and RegMean++ for merging. RegMean++ leverages representations from the merge model for merging, enabling accurate alignment with its behavior.
@@ 2 Background and Related Work
Model Soups (wortsman2022modelsoupsaveragingweights; choshen2022fusing), a well-known approach that performs merging by simply taking the average of all candidate models’ parameters, for enhancing distribution robustness (wortsman2022modelsoupsaveragingweights), creating merge models with multi-task or multi-modality capabilities (sung2023empirical; ilharco2023editingmodelstaskarithmetic; yadav2023tiesmergingresolvinginterferencemerging).
Task Arithmetic (ilharco2023editingmodelstaskarithmetic) introduces “task vector”, a new concept that quantifies the task-specific knowledge of a candidate model, by measuring the difference between the candidate model parameters and the base model parameters. However, yadav2023tiesmergingresolvinginterferencemerging point out that Task Arithmetic suffers from conflicts where different task vectors update the same parameters in opposite directions. They proposed TIES-Mering, a three-step method that first removes low-magnitude (noisy) parameters, then resolves sign conflicts, and finally aggregates only the non-conflicting parameters. DARE (yu2024language) proposes to pre-process each task vector independently by applying a Bernoulli mask to randomly zero out its parameters, effectively performing a dropout-like pruning before merging. TSV-M (gargiulo2025task) extracts task singular vectors via singular value decomposition (SVD) on layer-wise weight updates, selects top-$k$ components for a low-rank representation, and applies whitening to decorrelate subspaces before merging. marczakno argue that effective merging depends on how well merge updates span each task’s principal subspaces. They propose Iso-C, which sums task-specific updates, performs SVD, and replaces singular values with their average to create an isotropic spectrum. Iso-CTS, an Iso-C’s variant, adds top singular directions from each task’s residual, then orthogonalizes and isotropically scales the result.
Inspired by test-time adaptation schemes (wangtent; liang2025comprehensive), AdaMerging (yang2024adamergingadaptivemodelmerging) adaptively learns the merging coefficient for each layer of each task vector by minimizing the entropy on unlabeled test data, using it as a surrogate objective to refine the merge model’s performance across multiple tasks. DOGE AM (wei2025modeling) applies adaptive projective gradient descent at test time by jointly tuning a small modification vector and layer-wise merging coefficients to minimize prediction entropy on unlabeled inputs. DOGE AM projects updates orthogonally to a shared subspace to resolve task conflicts without hurting the shared knowledge.
Other approaches rely on data statistics, such as Fisher Merging (matena2022mergingmodelsfisherweightedaveraging) that requires computing Fisher information matrices. RegMean (jin2025datalessknowledgefusionmerging) gets rid of expensive gradient computation, formulates merging as a regression problem, then comes up with a computationally efficient and explainable closed-form solution.
@@
We denote matrices by boldface uppercase letters (e.g., $\bm{X}$, $\bm{W}$, $\bm{\Lambda}$), scalars by lowercase letters (e.g., $\alpha$, $l$). For operators, we denote $||\cdot||$ the Euclidean norm and $\text{tr}(\cdot)$ the trace of a matrix. Following  jin2025datalessknowledgefusionmerging, we consider the training-free merging framework. In such scenarios, we have access to a pool of multiple candidate models. Each candidate model is denoted as $f_{i}:\mathbb{R}^{N_{i}\times d}\to\mathbb{R}^{N_{i}\times|C|}$, and is fine-tuned on a task-specific dataset $\mathcal{D}_{i}=\{(\bm{X}_{i},\bm{Y}_{i})\}$. Here, $\bm{X}_{i}\in\mathbb{R}^{N_{i}\times d}$ is the batch-input of $N_{i}$ samples each of dimensionality $d$; and $\bm{Y}_{i}\in\mathbb{R}^{N_{i}\times|C|}$ is the corresponding target outputs, where $C$ is the set of classes. Our target is to find a merging function that takes a set of $K$ candidate models $f_{i}$, $i\in[1..K]$ and some data points, e.g., the task-specific training samples or held-out out-of-domain samples, as the inputs and returns a merge model $f_{{M}}$. We assume that the model architecture for all models is the same.
Regression Mean (RegMean;  (jin2025datalessknowledgefusionmerging))
@@
where $\bm{W}^{(l)}_{M}$ is the merge linear layer’s weights at the same position as $\bm{W}^{(l)}_{i}$ in the model $f_{i}$, $\bm{\Lambda}^{(l)}_{i}=\frac{1-\alpha}{\alpha}\text{diag}\left((\bm{X}^{(l)}_{i})^{\top}\bm{X}^{(l)}_{i}\right)\succeq 0$ is a regularization-strength diagonal matrix for $\bm{W}_{i}$, where $0\leq\alpha\leq 1$ is a predetermined scaling factor. Minimizing the loss $\mathcal{L}^{\text{RegMean}}$ in Eqn. 1 means finding $\bm{W}^{(l)}_{M}$ that ➀ approximates the behaviors of all candidate models while ➁ enforcing a regularization that keeps linear layer’s weights of the merge model $\bm{W}^{(l)}_{M}$ close to those of the candidate model $\bm{W}^{(l)}_{i}$. Minimizing $\mathcal{L}^{\text{RegMean}}$ describes a linear regression problem, where the inputs are $\bm{X}^{(l)}_{1},...,\bm{X}^{(l)}_{K}$] and the target outputs are $\bm{X}^{(l)}_{1}\bm{W}^{(l)}_{1},...,\bm{X}^{(l)}_{K}\bm{W}^{(l)}_{K}$] . This objective has a closed-form solution as:
@@
where $\widehat{\bm{G}}^{(l)}_{i}=\alpha\bm{G}^{(l)}_{i}+(1-\alpha)\text{diag}(\bm{G}^{(l)}_{i})=\alpha(\bm{X}^{(l)}_{i})^{\top}\bm{X}^{(l)}_{i}+(1-\alpha)\text{diag}((\bm{X}^{(l)}_{i})^{\top}\bm{X}^{(l)}_{i})$. Proof can be found in Appendix A. Other types of weights in the transformer layer are merged using averaging.
@@ 3.1 Motivation
Our first start is to revisit the underlying RegMean’s merging mechanism. RegMean operates by independently applying its closed-form solution to linear layers, including those within MLPs (up and down projections) and attention heads (key, query, and value matrices), across all candidate models. These components are well-known to store most of the model’s learned knowledge (meng2022locating), which may explain the effectiveness of RegMean.
Deep networks consist of multiple non-linear components, such as GELU and LayerNorm, which are interleaved with linear components. Due to the non-linear properties, even a small change in the input might potentially cause a large, unpredictable shift in the output. RegMean overlooks the information flow and feature transformations occurring at both the intra-layer level, i.e., within each layer, and cross-layer level of the merge model.
We hypothesize that incorporating those inference dynamics, that is, intra- and cross-layer dependencies, into RegMean’s merging objective is crucial for improving the merge model’s utilities and generalization. Inspired by this discussion, we introduce RegMean++, a simple yet powerful extension of RegMean in Section 3.2 below.
3.2 RegMean++ for Model Merging
@@
Let us fine-grain denote $\bm{X}^{(l,j)}_{i}$ be the input features of the $j$-th linear layer $\bm{W}^{(l,j)}_{i}$ in model $f_{i}$ at transformer layer $l$. RegMean computes the $j$-th merge linear layer weights $\bm{W}^{(l,j)}_{M}$ by Eqn. 2. In this closed-form solution, the merge weight $\bm{W}^{(l,j)}_{M}$ is determined by the individual weights and data statistics $\bm{G}^{(l,j)}_{i}=(\bm{X}_{i}^{(l,j)})^{\top}\bm{X}_{i}^{(l,j)}$, which captures the dependencies among input features across all candidate models.
Algorithm. Similar to RegMean, given a $j$-th linear layer at the transformer layer $l$, RegMean++ computes the inner product matrix as $\bm{G}_{i}^{(l,j)}=(\bm{X}^{(l,j)}_{i})^{\top}\bm{X}^{(l,j)}_{i}$. The key difference between RegMean++ and RegMean lies in how input feature $\bm{X}^{(l,j)}_{i}$ is obtained: For input features that are activations (cushion representations between transformer layers), RegMean++ computes $\bm{X}^{(l,j)}_{i}$ based on the activations produced by the previous merge layer $f_{M}^{(l-1)}$ in the merge model, that is, $\bm{X}^{(l)}_{i}=f_{M}^{(l-1)}(\bm{X}^{(l-1)}_{i})$ while RegMean relies on the activations produced by the previous candidate layer $f_{i}^{(l-1)}$ in the candidate model, that is, $\bm{X}^{(l)}_{i}=f_{i}^{(l-1)}(\bm{X}^{(l-1)}_{i})$. Similar to RegMean, all other parameters in the transformer layer, such as embeddings and biases, are merged via simple averaging. RegMean++ inherits RegMean’s advantages, but introduces a trade-off between model performance and computational cost. During the statistic‐collection phase, RegMean++ incurs additional forward passes, which are computed in the merge model to collect the inner-product matrices, yet the merging time equals that of RegMean. Our RegMean++ pseudocode is described in Algorithm 1. Comparison between RegMean and RegMean++ is described in Figure 1.
@@ 4 Experiments
4.1 Models and Datasets
Vision classification tasks.
Following prior works (ilharco2023editingmodelstaskarithmetic; yang2024adamergingadaptivemodelmerging; wei2025modeling), we evaluate the effectiveness of multi-task model merging methods on eight standard datasets including: SUN397 (xiao2016sun), Stanford Cars (Cars) (krause20133d), RESISC45 (cheng2017remote), EuroSAT (helber2019eurosat), SVHN (netzer2011reading), GTSRB (stallkamp2011german), MNIST (lecun1998gradient), and DTD (cimpoi2014describing). Following wang2024localizing, we also employ $12$ additional tasks for evaluation on the sustainability.
We assess the performance of merging methods on three CLIP model variants (radford2021learning) with ViT-B/32, ViT-B/16, and ViT-L/14. For the candidate models, we employ off-the-shelf checkpoints from tang2024fusionbench.
Language generation tasks.
Following he2025mergebench, we evaluate the merge models on $11$ datasets reflecting five domains: (1) instruction following: IFEval (zhou2023instruction), (2) mathematics: GSM8K (cobbe2021training), (3) multilingual understanding (on French, Spanish, German, and Russian): Multilingual MMLU, Multilingual ARC, and Multilingual Hellaswag (lai-etal-2023-okapi), (4) coding: HumanEval+ and MBPP+ (liu2023your), and (5) safety: WildGuardTest (han2024wildguard), HarmBench (mazeika2024harmbench), DoAnythingNow (shen2024anything), and XSTest (rottger2024xstest).
We assess the performance of merging methods on two Llama 3 variants (grattafiori2024llama3herdmodels) with Llama-3.2-3B and Llama-3.1-8B. For the candidate models, we employ off-the-shelf checkpoints from he2025mergebench.
Details of these datasets’ description and models can be found in Appendix B.1 and Appendix B.2.
4.2 Comparison Methods
We compare RegMean++ against $11$ recent advanced model merging methods spanning three categories: (1) Data-Free methods: Model Soups (wortsman2022modelsoupsaveragingweights), Task Arithmetic (ilharco2023editingmodelstaskarithmetic), TIES-Merging (yadav2023tiesmergingresolvinginterferencemerging), TSV-M (gargiulo2025task), DOGE TA (wei2025modeling), Iso-C and Iso-CTS (marczakno). (2) Training-Free methods: Fisher Merging (matena2022mergingmodelsfisherweightedaveraging) and RegMean (jin2025datalessknowledgefusionmerging). (3) Test-Time Adaptation: AdaMerging (yang2024adamergingadaptivemodelmerging) and DOGE AM (wei2025modeling). In addition, we also consider the MTL as an upper-bound performance. Details can be found in Appendix B.3.
4.3 Experimental Setup
We employ FusionBench (tang2024fusionbench) and MergeBench (he2025mergebench) for merging evaluation on vision and language tasks, respectively. For vision tasks, we report the accuracy and normalized accuracy (See Appendix C.1 for details). For language tasks, multiple task-specific metrics are employed (See Appendix C.2 for details). All of the metrics are calculated on the tasks’ test split. Hyperparameters for settings are specified in their respective subsections, and further detailed in Appendix C.3. Due to space constraints, we present key results in the main text and defer full experimental setups and additional results to Appendix C.4 and Appendix D, respectively. Implementation and guidelines for reproducing those results are attached to the supplemental materials.